The heat of combustion of octane, C_8H_18, can be
Answer #1

heat released by 1.02 g Octane = mw*Cw*delta T
= 1000 g * 4.184 J/goC * (35.3-22) oC
= 55647.2 J

number of moles of Octane = mass/molar mass
= 1.02 g/ (8*12 + 18*1) g/mol
= 8.95*10^-3 mol

H = -55647.2 / (8.95*10^-3)
= -6.219*10^6 J/mol
= -6219 KJ/mol

Answer: -6219 KJ/mol